OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of −3 and passes through the point (−2,1).

OpenStudy (anonymous):

I know that it would be Y= -3x I just don't understand how to solve the other part of the question...

OpenStudy (michele_laino):

please you have to apply the formula below: \[y-y _{1}=m(x-x _{1})\] where m is your slope, and (x_1,y_1) is your point

OpenStudy (texaschic101):

not necessarily...you could just use y = mx + b

OpenStudy (anonymous):

\[Y = -3x - 5 ?\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (michele_laino):

Thank you!!

OpenStudy (yamyam70):

@Michele_Laino @texaschic101 , where did he get the value of b ? Im confused

OpenStudy (texaschic101):

which formula do you want to use...because you can either use point slope (y - y1 = m(x - x1) or you can use slope intercept (y = mx + b) ??

OpenStudy (michele_laino):

@yamyam70 please substitute x=-2 and y=1, and m=-3, in that formula above

OpenStudy (yamyam70):

lets go with , y = mx + b @Michele_Laino oh yeah, thanks! :D

OpenStudy (michele_laino):

@yamyam70 Thanks!

OpenStudy (yamyam70):

@texaschic101 is much simpler y= mx +b :) thank you too :)

OpenStudy (texaschic101):

I think so to, but either way will work

OpenStudy (yamyam70):

after all, every formula is just derived from one another

OpenStudy (texaschic101):

so true

OpenStudy (yamyam70):

alrighty :)